### v3

@@ 5 Abelian Group Isomorphism
Chattopadhyay, Torán, and Wagner [CTW13] established a $\textsf{TC}^{0}(\mathsf{FOLL})$ upper bound on this problem. Grochow & Levet [GL23, Theorem 5] gave a tighter analysis of their algorithm, placing it in the sub-class $\forall^{\log n}\mathsf{MAC^{0}}(\mathsf{FOLL})$. [^9] Grochow & Levet consider $\forall^{\log n}\mathsf{MAC^{0}}\circ\mathsf{FOLL}$, where $\circ$ denotes composition (see [GL23] for a precise formulation). We note that as $\mathsf{AC^{0}}\circ\mathsf{FOLL}=\mathsf{FOLL}=\mathsf{AC^{0}}(\mathsf{FOLL})$, we have that $\forall^{\log n}\mathsf{MAC^{0}}\circ\mathsf{FOLL}=\forall^{\log n}\mathsf{MAC^{0}}(\mathsf{FOLL})$. Thus, Thm. 5.1 improves upon the previous bound of $\forall^{\log n}\mathsf{MAC^{0}}(\mathsf{FOLL})$ obtained by Grochow & Levet. We note that Chattopadhyay, Torán, & Wagner also established an upper bound of $\mathsf{L}$ for this problem, which is incomparable to the result of Grochow & Levet (ibid.). We improve upon both these bounds by (i) showing that $O(\log\log n)$ non-deterministic bits suffice instead of $O(\log n)$ bits, and (ii) using an $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ circuit for order finding rather than an $\mathsf{FOLL}$ circuit. We note that while $\forall^{\log n}\mathsf{MAC^{0}}(\mathsf{FOLL})$ is contained in $\mathsf{TC^{0}}(\mathsf{FOLL})$, it is open whether this containment is strict. In contrast, Cor. 5.3 shows that our new bound of $\forall^{\log\log n}\mathsf{MAC}^{0}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ is a class that is in fact strictly contained in $\mathsf{L}\cap\textsf{TC}^{0}(\mathsf{FOLL})$.
Proof of Thm. 5.1.
@@
However, to identify elements of order $p^{e}$, we will need the number $p^{e}$ explicitly, in its full binary expansion. First we show that we can get $p^{e}$ explicitly if we can get $p$ explicitly. Once we have $p$ in its binary expansion, the function $(p,e)\mapsto p^{e}$ is a function of $O(\log n)$ bits, so can be computed by an $\mathsf{AC}^{0}$ circuit of size $2^{O(\log n)}=n^{O(1)}$ (see Obs. 2.2). Thus all that remains is to get $p$ explicitly.
@@ 6 Isomorphism for Groups of Almost All Orders
Fix a group $G$ of order $n$. We will first discuss how to decompose $G=H\ltimes B$, as prescribed by [DW22, Theorem 2.5]. Without loss of generality, suppose that $p_{1},\ldots,p_{\ell}$ ($\ell\leq k$) are the unique primes where $p_{i}$ ($i\in[\ell]$) divides $n$ only once and $p_{i}>\log\log n$. Now as each $p_{i}$ can be represented as a string of length $\leq\lceil\log(n)\rceil$, we may in $\mathsf{AC}^{0}$ compute $\overline{p}:=p_{1}\cdots p_{\ell}$ (Obs. 2.2). Using Lem. 3.1, we may in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ identify an element $g\in G$ of order $\overline{p}$.
@@ 7.1 MGS for Groups in $\mathsf{AC}^{1}(\mathsf{L})$
  • Case 1: Suppose that $N=N_{i}/N_{i-1}$ is Abelian. By [LM94, Theorem 4], we have three cases:
    • Case 1a: $G/N_{i-1}=\langle g_{1},\cdots,g_{d}\rangle$. We may test whether this holds in $\mathsf{L}$ using a membership test.
    • Case 1b: We have $G/N_{i-1}=\langle g_{1},\cdots,g_{i},g_{j}n,g_{j+1},\cdots,g_{d}\rangle$ for some $j\in[d]$ and some $n\in N$. There are at most $d\cdot(|N|-1)$ generating sets to consider in this case and we can test each of them in $\mathsf{L}$.
    • Case 1c: If neither Cases 1a or 1b hold, then we necessarily have that $G/N_{i-1}=\langle g_{1},\cdots,g_{d},x\rangle$ for some any non-identity element $x\in N$.
    Note that there are at most $d\cdot|N|$ generating sets to consider, we may construct a minimum generating set for $G/N_{i-1}$ in $\mathsf{L}$ using a membership test.
  • Case 2: Suppose instead that $N=N_{i}/N_{i-1}$ is non-Abelian. Then by [LT24, Corollary 13] we have the following cases:
    • Case 2a: $G/N_{i-1}=\langle g_{1},\cdots,g_{d}\rangle$. We may test whether this holds in $\mathsf{L}$ using a membership test.
    • Case 2b: If Case 2a does not hold, then we have the following case. Let $\eta_{G}(N)$ denote the the number of factors in a chief series with order $\lvert N\rvert$. Let $u=\max\{d,2\}$ and $t=\min\{u,\lceil\frac{8}{5}+\log_{\lvert N\rvert}\eta_{G}(N)\rceil\}$. Then there exist $n_{1},\ldots,n_{t}\in N_{i-1}$ such that $G/N_{i-1}=\langle g_{1}n_{1},\cdots,g_{t}n_{t},g_{t+1},\cdots,g_{d}\rangle$.
      By [LT24, Corollary 13], there are at most $\lvert N\rvert^{\lceil\frac{8}{5}+\log_{\lvert N\rvert}\eta_{G}(N)\rceil}$ generating sets of this form. We may write down these generating sets in parallel with a single $\mathsf{AC}^{0}$ circuit and test whether each generate $G/N_{i-1}$ in $\mathsf{L}$ using a membership test.

### v4

@@ 4 Abelian Group Isomorphism
Chattopadhyay, Torán, and Wagner [CTW13] established a $\textsf{TC}^{0}(\mathsf{FOLL})$ upper bound on this problem. Grochow & Levet [GL23, Theorem 5] gave a tighter analysis of their algorithm, placing it in the sub-class $\forall^{\log n}\mathsf{MAC^{0}}(\mathsf{FOLL})$. [^15] Grochow & Levet consider $\forall^{\log n}\mathsf{MAC^{0}}\circ\mathsf{FOLL}$, where $\circ$ denotes composition (see [GL23] for a precise formulation). We note that as $\mathsf{AC^{0}}\circ\mathsf{FOLL}=\mathsf{FOLL}=\mathsf{AC^{0}}(\mathsf{FOLL})$, we have that $\forall^{\log n}\mathsf{MAC^{0}}\circ\mathsf{FOLL}=\forall^{\log n}\mathsf{MAC^{0}}(\mathsf{FOLL})$. Thus, Thm. 4.1 improves upon the previous bound of $\forall^{\log n}\mathsf{MAC^{0}}(\mathsf{FOLL})$ obtained by Grochow & Levet. We note that Chattopadhyay, Torán, & Wagner also established an upper bound of $\mathsf{L}$ for this problem, which is incomparable to the result of Grochow & Levet (ibid.). We improve upon both these bounds by (i) showing that $O(\log\log n)$ non-deterministic bits suffice instead of $O(\log n)$ bits, and (ii) using an $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ circuit for order finding rather than an $\mathsf{FOLL}$ circuit. We note that while $\forall^{\log n}\mathsf{MAC^{0}}(\mathsf{FOLL})$ is contained in $\mathsf{TC^{0}}(\mathsf{FOLL})$, it is open whether this containment is strict. In contrast, Cor. 4.3 shows that our new bound of $\forall^{\log\log n}\mathsf{MAC}^{0}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ is a class that is in fact strictly contained in $\mathsf{L}\cap\textsf{TC}^{0}(\mathsf{FOLL})$.
Proof of Thm. 4.1.
@@
However, to identify elements of order $p^{e}$, we will need the number $p^{e}$ explicitly, in its full binary expansion. First we show that we can get $p^{e}$ explicitly if we can get $p$ explicitly. Once we have $p$ in its binary expansion, the function $(p,e)\mapsto p^{e}$ is a function of $O(\log n)$ bits, so can be computed by an $\mathsf{AC}^{0}$ circuit of size $2^{O(\log n)}=n^{O(1)}$ (see Thm. 2.2). Thus all that remains is to get $p$ explicitly.
@@ 5 Isomorphism for Groups of Almost All Orders
Fix a group $G$ of order $n$. We will first discuss how to decompose $G=H\ltimes B$, as prescribed by [DW22, Theorem 2.5]. Without loss of generality, suppose that $p_{1},\ldots,p_{\ell}$ ($\ell\leq k$) are the unique primes where $p_{i}$ ($i\in[\ell]$) divides $n$ only once and $p_{i}>\log\log n$. Now as each $p_{i}$ can be represented as a string of length $\leq\lceil\log(n)\rceil$, we may in $\mathsf{AC}^{0}$ compute $\overline{p}:=p_{1}\cdots p_{\ell}$ (Thm. 2.2). Using Lem. 3.1, we may in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ identify an element $g\in G$ of order $\overline{p}$.
@@
6 Quasigroup Isomorphism
In this section, we establish the following.
Theorem 6.1.
Quasigroup Isomorphism belongs to $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$. In particular, it can be solved by $\mathsf{quasiAC^{0}}$ circuits of depth 4.
Proof of Thm. 6.1.
Let $G$ and $H$ be quasigroups given as their multiplication tables. We assume that the elements of the quasigroups are indexed by integers $1,\dots,\left|\mathinner{G}\right|$. If $\left|\mathinner{G}\right|\neq\left|\mathinner{H}\right|$ (this can be tested in $\mathsf{DTIME}(\log n)$ by a standard binary search), we know that $G$ and $H$ are not isomorphic. Otherwise, let us write $n=\left|\mathinner{G}\right|$ and $k=\left\lceil\mathinner{2\log n}\right\rceil+1$ (while [CTW13, Theorem 3.3] only states that some $k\in O(\log n)$ suffices, the corresponding proof states that for $k=\left\lceil\mathinner{2\log n}\right\rceil+1$, there is actually a cube generating set).
  • that these sequences are actually cube generating sequences,
  • that $g_{i}\mapsto h_{i}$ induces an isomorphism.
Let us describe the first point for $G$ (for $H$ this follows exactly the same way): we universally verify for every element $g\in G$ (which can be encoded using $O(\log n)$ bits, hence, $\forall^{\log n}$) that we can existentially guess a sequence $(e_{1},\dots,e_{k})\in\{0,1\}^{k}$ (i.e. $\exists^{\log n}$) such that $g=P(g_{0}g_{1}^{e_{1}}\cdots g_{k}^{e_{k}})$. We can compute this product in $\mathsf{DTISP}(\log^{2+o(1)}n,\log n)$ by multiplying from left to right. [^16] Here, we could impose the additional requirement that the length of each row/column of the multiplication table is padded up to a power of two in order to get a bound of $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\log^{2}n,\log n)$) Each multiplication can be done in time $O(\log^{1+o(1)}n)$ because we simply need to compute $i+j\cdot n$ for two addresses $i,j$ of quasigroup elements, write the result on the index tape and then read the corresponding product of group elements from the multiplication table. Moreover, note that for this procedure we only need to store one intermediate result on the working tape at any time. Thus, computing the product $P(g_{0}g_{1}^{e_{1}}\cdots g_{k}^{e_{k}})$ can be done in time $O(\log^{2+o(1)}n)$ and space $O(\log n)$ and it can be checked whether the result is $g$.
Again the products can be computed in time $O(\operatorname{poly}\log n)$ and space $O(\log n)$ as outlined above. Now, it remains to observe that we can combine the two $\forall^{\log n}$ blocks into one $\forall^{\log n}$ block because the check whether $g_{i}\mapsto h_{i}$ induces an isomorphism does not depend on the existentially guessed bits in the first check.
The depth 4 circuit bound then follows from the DNF for $\mathsf{DTIME}(\text{polylog}(n))$ (Fact 2.4): the quantifier blocks $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}$ turn into $\bigvee^{\log^{2}n}\bigwedge^{\log n}\bigvee^{\log n}$, and then the disjunction at the top of the DNF for $\mathsf{DTIME}(\text{polylog}(n))$ can be merged into the final $\bigvee$ from the quantifier blocks. ∎
Corollary 6.2.
The following problems belong to $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
  1. (a)
    Latin Square Isotopy. In particular, this problem belongs to $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$, which yields depth $4$ $\mathsf{quasiAC^{0}}$ circuits.
  2. (b)
    Isomorphism testing of two Steiner triple systems. In particular, this problem belongs to
    $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$, which yields depth $4$ $\mathsf{quasiAC^{0}}$ circuits.
  3. (c)
    Isomorphism testing of Latin square graphs.
  4. (d)
    Isomorphism testing of Steiner $(t,t+1)$-designs.
  5. (e)
    Isomorphism testing of pseudo-STS graphs.
For Latin Square Isotopy and isomorphism testing of Steiner triple systems, a careful analysis yields depth-$4$ $\mathsf{quasiAC^{0}}$ circuits. In contrast, the reductions from [Lev23] for isomorphism testing of Latin square graphs, Steiner $(t,t+1)$-designs, and pseudo-STS graphs all use circuits of depth at least $4$. Obtaining reductions of depth less than $4$ will likely require new techniques.
Proof.
We proceed as follows.
  1. (a)
    We note that the reductions outlined in [Mil78, Theorem 2] and [Lev23, Remark 1.6] in fact allow us to determine whether two quasigroup are isotopic, and not just main class isomorphic. Thus, we have an $\mathsf{AC^{0}}$-computable disjunctive truth-table reduction from Latin Square Isotopy to Quasigroup Isomorphism. This suffices to yield the bound of $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
    We now turn to establishing the stronger $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ bound. We carefully analyze the $\exists^{\log^{2}n}\mathsf{L}\cap\exists^{\log^{2}n}\mathsf{FOLL}$ algorithm from [Lev23, Section 3] using cube generating sequences. Our analysis adapts the proof technique of Theorem 6.1. Let $Q_{1},Q_{2}$ be quasigroups, and let $k\in O(\log n)$. We begin by guessing three cube generating sequences $\bar{a}=(a_{0},a_{1},\ldots,a_{k}),\bar{b}=(b_{0},b_{1},\ldots,b_{k}),\bar{c}=(c_{0},c_{1},\ldots,c_{k})$ for $Q_{1}$, and $\bar{a^{\prime}}=(a_{0}^{\prime},a_{1}^{\prime},\ldots,a_{k}^{\prime}),\bar{b^{\prime}}=(b_{0}^{\prime},b_{1}^{\prime},\ldots,b_{k}^{\prime}),\bar{c^{\prime}}=(c_{0}^{\prime},c_{1}^{\prime},\ldots,c_{k}^{\prime})$ for $Q_{2}$. This requires $O(\log^{2}n)$ non-deterministic bits ($\exists^{\log^{2}n}$). From the proof of Theorem 6.1, we may check in $\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ whether each of $\bar{a},\bar{b},\bar{c}$ generates $Q_{1}$, and whether $\bar{a^{\prime}},\bar{b^{\prime}},\bar{c^{\prime}}$ generates $Q_{2}$.
    So now suppose that each of $\bar{a},\bar{b},\bar{c}$ generates $Q_{1}$, and each of $\bar{a^{\prime}},\bar{b^{\prime}},\bar{c^{\prime}}$ generates $Q_{2}$. Let $\alpha,\beta,\gamma:Q_{1}\to Q_{2}$ be the bijections induced by the respective maps $\bar{a}\mapsto\bar{a^{\prime}},\bar{b}\mapsto\bar{b^{\prime}}$, and $\bar{c}\mapsto\bar{c^{\prime}}$. Now for each pair $(g,h)\in Q_{1}$, there exist $x,y,z\in\{0,1\}^{k}$ s.t.:
    $g=a_{0}a_{1}^{x_{1}}\cdots a_{k}^{x_{k}},$
    $h=b_{0}b_{1}^{y_{1}}\cdots b_{k}^{y_{k}}$
    $gh=c_{0}c_{1}^{z_{1}}\cdots c_{k}^{z_{k}}.$
    We require $O(\log n)$ universally quantified co-nondeterministic bits ($\forall^{\log n}$) to represent $(g,h)$, and $O(\log n)$ existentially quantified non-deterministic bits to represent $x,y,z$ ($\exists^{\log n}$). Given $x,y,z$, we may write down $g,h,gh$ in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ (following the proof of Theorem 6.1). Similarly, let:
    $g^{\prime}=a_{0}^{\prime}(a_{1}^{\prime})^{x_{1}}\cdots(a_{k}^{\prime})^{x_{k}},$
    $h^{\prime}=b_{0}^{\prime}(b_{1}^{\prime})^{y_{1}}\cdots(b_{k}^{\prime})^{y_{k}}$
    $\ell^{\prime}=c_{0}(c_{1}^{\prime})^{z_{1}}\cdots(c_{k}^{\prime})^{z_{k}}.$
    We can check in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ whether $g^{\prime}h^{\prime}=\ell^{\prime}$. Similar to the proof of Theorem 6.1, we can combine the two $\forall^{\log n}$ blocks into one $\forall^{\log n}$ block, as the check whether $(\bar{a},\bar{b},\bar{c})\mapsto(\bar{a^{\prime}},\bar{b^{\prime}},\bar{c^{\prime}})$ preserves the quasigroup operation (the homotopism condition) is independent of whether each of $\bar{a},\bar{b},\bar{c}$ generates $Q_{1}$, and each of $\bar{a^{\prime}},\bar{b^{\prime}},\bar{c^{\prime}}$ generates $Q_{2}$.
    By Fact 2.4, an $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ machine can be simulated by a depth-$2$ $\mathsf{quasiAC^{0}}$ circuit, taking care of the final $\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$. The additional quantifiers $\exists^{\log^{2}n}\forall^{\log n}$ then yield depth-$4$ $\mathsf{quasiAC^{0}}$ circuits.
  2. (b)
    Given a Steiner triple system, we may obtain a quasigroup $Q$ in the following manner. For each block $\{a,b,c\}$ in the Steiner triple system, we include the products $ab=c,ba=c,ac=b,ca=b,bc=a,cb=a$. We can write down the multiplication table using an $\mathsf{AC^{0}}$ circuit, which suffices to obtain a bound of $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
    However, we can use the blocks of the Steiner triple system to look up the relevant products in $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$, and so we need not write down the multiplication table. This allows us to directly apply the proof of Theorem 6.1, which yields a bound of
    $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n)),$
    and so we obtain depth-$4$ $\mathsf{quasiAC^{0}}$ circuits, as desired. (Analyzing the $\mathsf{AC^{0}}$ reduction in terms of circuits, yielded only depth 6 in the end, whereas realizing the reduction can be computed in $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ allows us to get depth 4 overall.)
  3. (c)
    Given two Latin square graphs $G_{1},G_{2}$, we may recover corresponding Latin squares $L_{1},L_{2}$ in $\mathsf{AC^{0}}$ (cf. [Lev23, Lemma 3.9]). Now $G_{1}\cong G_{2}$ if and only if $L_{1}$ and $L_{2}$ are main class isomorphic [Mil78, Lemma 3]. We may decide whether $L_{1},L_{2}$ are main class isomorphic using an $\mathsf{AC^{0}}$-computable disjunctive truth-table reduction to Quasigroup Isomorphism (cf. [Lev23, Remark 1.6]). By Fact 2.7, $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ is closed under $\mathsf{AC^{0}}$-computable dtt reductions, thus yielding the bound of $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
  4. (d)
    This immediately follows from the fact that isomorphism testing of Steiner $(t,t+1)$-designs is $\exists^{\log^{2}n}\mathsf{AC^{0}}$-reducible to isomorphism testing of Steiner triple systems [BW13] (cf., [Lev23, Corollary 4.11]).
  5. (e)
    Bose [Bos63] previously showed that pseudo-STS graphs with $>67$ vertices are STS graphs. Now given a block-incidence graph from a Steiner $2$-design with bounded block size, we can recover the underlying design in $\mathsf{AC^{0}}$ [Lev23, Proposition 4.7]. This yields the desired bound.
Remark 6.3.
@@ 7.1 MGS for Groups in $\mathsf{AC}^{1}(\mathsf{L})$
  • Case 1: Suppose that $N=N_{i}/N_{i-1}$ is Abelian. By [LM94, Theorem 4], we have two cases:
    • Case 1a: We have $G/N_{i-1}=\langle g_{1},\cdots,g_{i},g_{j}n,g_{j+1},\cdots,g_{d}\rangle$ for some $j\in[d]$ and some $n\in N$ (possibly $n=1$). There are at most $d\cdot|N|$ generating sets to consider in this case and we can test each of them in $\mathsf{L}$.
    • Case 1b: If Case 1a does not hold, then we necessarily have that $G/N_{i-1}=\langle g_{1},\cdots,g_{d},x\rangle$ for some any non-identity element $x\in N$.
    Note that there are at most $d\cdot|N|+1$ generating sets to consider, we may construct a minimum generating set for $G/N_{i-1}$ in $\mathsf{L}$ using a membership test.
  • Case 2: Suppose instead that $N=N_{i}/N_{i-1}$ is non-Abelian. Then by [LT24, Corollary 13], the following holds. Let $\eta(G)$ denote the number of factors in a chief series with order $|N|$. Let $u=\max\{d,2\}$ and $t=\min\{u,\lceil\frac{8}{5}+\log_{\lvert N\rvert}\eta_{G}(N)\rceil\}$. Then there exist $n_{1},\ldots,n_{t}\in N_{i-1}$ (possibly $n_{1}=\cdots=n_{t}=1$) such that $G/N_{i-1}=\langle g_{1}n_{1},\cdots,g_{t}n_{t},g_{t+1},\cdots,g_{d}\rangle$.
    By [LT24, Corollary 13], there are at most $\lvert N\rvert^{\lceil\frac{8}{5}+\log_{\lvert N\rvert}\eta_{G}(N)\rceil}$ generating sets of this form. As $\log_{|N|}\eta_{G}(N)\in O(1)$, we may write down these generating sets in parallel with a single $\mathsf{AC}^{0}$ circuit and test whether each generate $G/N_{i-1}$ in $\mathsf{L}$ using Membership.